Find the slope of the line passing through the points (5,7) and (-4,2).

Found 2 solutions by stanbon, Maths68:
Answer by stanbon(75887)

Find the slope of the line passing through the points (5,7) and (-4,2).
----
m = (7-2)/(5--4) = 5/9
==========================
Cheers,
Stan H.

Answer by Maths68(1474)

Solution
Points (5,7) and (-4,2)
Put the values of two points in the equation of straight line point-slope form
m=(y2-y1)/(x1-x2)
m=(2-7)/(-4-5)
m=-5/-9
m=5/9
Slope of the line = m = 5/9